# Characteristic Polynomials and Hypergraph Generating Functions via Heaps of Pieces

### v1

@@ Abstract
It is a classical result due to Jacobi in algebraic combinatorics that the generating function of closed walks at a vertex $u$ in a graph $\mathcal{G}$ is determined by the rational function
$\frac{\phi_{\mathcal{G}-u}(t)}{\phi_{\mathcal{G}}(t)}$
where $\phi_{\mathcal{G}}(t)$ is the characteristic polynomial of $\mathcal{G}$. In this paper, we show that the corresponding rational function for a hypergraph is also a generating function for some combinatorial objects in the hypergraph.
We make use of the Heaps of Pieces framework, developed by Viennot, demonstrating its use on graphs, digraphs, and multigraphs before using it on hypergraphs. In the case of a graph $\mathcal{G}$, the pieces are cycles and the concurrence relation is sharing a vertex. The pyramids with maximal piece containing a vertex $u\in V(\mathcal{G})$ are in one-to-one correspondence with closed walks at $u$. In the case of a hypergraph $\mathcal{H}$, connected “infragraphs” can be defined as the set of pieces, with the same concurrence relation: sharing a vertex. Our main results are established by analyzing multivariate resultants of polynomial systems associated to adjacency hypermatrices.
@@ 1 Introduction
Given a simple graph $\mathcal{G}$, let $\phi_{\mathcal{G}}(t)$ be the characteristic polynomial of its adjacency matrix $\mathbb{A}_{\mathcal{G}}$. For a vertex $u\in V(\mathcal{G})$, a well-known classical identity ([12, Equation (1), p. 52]) is,
$\sum_{d\geq 0}w_{d}^{u}t^{d}=\frac{t^{-1}\phi_{\mathcal{G}-u}(t^{-1})}{\phi_{\mathcal{G}}(t^{-1})}$ (1.1)
where the $d$-th coefficient $w_{d}^{u}$ is the number of closed walks of length $d$ at $u$, and $\mathcal{G}-u$ is the vertex-deleted subgraph of $\mathcal{G}$, obtained by deleting the vertex $u$, together with all the edges incident with $u$. This result is attributed to Jacobi in 1833 [16], see also [12, 11]. Note that the rational function on the right-hand side is equal to the generating function on the left-hand side, with non-negative integer coefficients. Thus, the function in question is an example of a broader class of functions of interest, the rational generating functions, investigated in [5].
@@
The theory of Heaps of Pieces is developed by Viennot (see [28, p. 36], [26, 3, 5, 18]). In the case of a finite simple graph $\mathcal{G}$, one takes cycles of $\mathcal{G}$ as pieces, and sharing a vertex as the concurrence relation. In Section 2.4, this approach is demonstrated and and laid out in detail, both for completeness and for reference. In particular, Corollary 2.7 establishes the bijection between closed walks of $\mathcal{G}$ at a fixed vertex $u$, and the pyramids with maximal piece containing $u$.
@@
When two infragraphs are disconnected, we show that the weight of the disjoint union is the product of the weights of the components, so the multiplicativity condition is satisfied. This setting generalizes the Heaps of Pieces framework outlined above for graphs of rank $2$ in the following sense: The weight of a connected infragraph $X$ of a graph $\mathcal{G}$ of rank $2$ is non-zero only if $X$ is indecomposable, i.e., it is a cycle or an edge (see [8]).
@@ 2.2 Operations on the Trails of an Eulerian Digraph
We denote by $E(\mathbf{w})=\{e_{1},\ldots,e_{d}\}$ the set of edges and by $V(\mathbf{w})=\{v_{0},\ldots,v_{d}\}$ the set of vertices of a walk $\mathbf{w}$. A walk $\mathbf{w}$ is called a trail, if its edges are distinct. A closed trail $\mathbf{w}$ of $X$ is an Eulerian trail when $E(\mathbf{w})=E(X)$. A path is a trail without repeated vertices. A closed trail without repeated internal vertices is said to be a simple closed trail. We denote by $\mathcal{T}(X)$ and $\mathcal{W}(X)$ the sets of closed trails and Eulerian (closed) trails of $X$, respectively. Given a vertex $u\in V(X)$, we write $\mathcal{T}^{u}(X)$ for the set of closed trails of $X$ at $u$ and $\mathcal{W}^{u}(X)$ for the set of Eulerian trails of $X$ at $u$. Given an edge $e\in E(X)$, we denote by $\mathcal{T}^{e}(X)$, the set of closed trails of $X$ ending at $e$ and by $\mathcal{W}^{e}(X)$, the set of Eulerian trails of $X$ ending at $e$.
@@ 2.4 Bijections between Pyramids and Trails/Walks
where, $v_{m}=v_{0}^{\prime}$ and $0\leq m\leq r$. Then, the concatenation of $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, denoted $\mathbf{w}_{1}\mathbf{w}_{2}$, is the trail:
@@
Trails of a multi-graph do not visit the same edge more than once. Given a simple graph $\mathcal{G}$, we have a bijection (cf. Corollary 2.7 below.) featuring walks in which the same edge may be visited more than once and pyramids made of not necessarily distinct pieces. Given a sequence of vertices of a simple graph $\mathcal{G}$, there is at most one walk visiting the sequence of vertices in the given order. This means that a walk is uniquely determined by its sequence of vertices.
Let $\mathcal{G}$ be a connected simple graph of rank $2$. As defined in Section 2.1, the set of cycles of $\mathcal{G}$ is denoted $\mathcal{B}(\mathcal{G})$. Since the graph $\mathcal{G}$ is simple, it can not have a digon (a cycle of length $2$). However, walks of a graph of length $2$ up to cyclic permutation of edges, called edgegons, are included in the set of pieces of $\mathcal{G}$: Let $\mathcal{Q}(\mathcal{G}):=\{[(v_{0},e,v_{1},e,v_{0})]_{\rho}:e=\{v_{0},v_{1}\}\in E(\mathcal{G})\}$. Write $\widetilde{B}(\mathcal{G}):=\mathcal{B}(\mathcal{G})\sqcup\mathcal{Q}(\mathcal{G})$ for the union of the set of cycles (of length $\geq 3$) and the set of edgegons of $\mathcal{G}$. (We have $[(v_{0},e,v_{1},e,v_{0})]_{\rho}=[(v_{1},e,v_{0},e,v_{1})]_{\rho}$, for each edge $e=\{v_{0},v_{1}\}\in E(\mathcal{G})$ and necessarily, a graph $\mathcal{G}$ has $|E(\mathcal{G})|$ many edgegons.)
@@
    Consider the simple graph $\mathcal{G}$, obtained by appending an edge to a triangle. We list the pieces of $\mathcal{G}$ below.
@@
By Veblen’s Theorem ([25, p. 87]), a connected simple graph $\mathcal{G}$ has an Eulerian trail, i.e., $\mathcal{W}(\mathcal{G})\neq\emptyset$ if and only if it is obtained by an edge-disjoint union of cycles of length $\geq 3$.
Next, we derive a bijection involving walks of a simple graph $\mathcal{G}$ and pyramids made of cycles and edgegons.
@@
Let $\mathcal{G}$ be a simple connected graph, $u\in V(\mathcal{G})$ and $e\in E(\mathcal{G})$. Define the mapping
$q:\mathcal{U}(\mathcal{G})\rightarrow\mathbbm{p}(\mathcal{G})$
@@
    $q\big{|}_{\mathcal{U}^{e}(\mathcal{G})}:\mathcal{U}^{e}(\mathcal{G})\rightarrow\mathbbm{p}^{e}(\widetilde{\mathcal{B}}(\mathcal{G}))$ is a bijection between closed walks of $\mathcal{G}$ ending at $e$ and pyramids of $\mathcal{G}$, with maximal piece containing $e$.
@@
    $q\big{|}_{\mathcal{U}^{u}(\mathcal{G})}:\mathcal{U}^{u}(\mathcal{G})\rightarrow\mathbbm{p}^{u}(\widetilde{\mathcal{B}}(\mathcal{G}))$ is a bijection between closed walks of $\mathcal{G}$ at $u$ and pyramids of $\mathcal{G}$, with maximal piece containing $u$.
@@ 2.5 The Characteristic Polynomial of a Graph as a sum over Trivial Heaps
We intend to apply Theorem 2.8 to the characteristic polynomial of a simple graph $\mathcal{G}$ of order $n$ and rank $2$, which is defined as the determinant $\text{det}(t\cdot I_{n}-\mathbb{A}_{\mathcal{G}})$, where $I_{n}$ is the identity matrix of dimension $n$ and $\mathbb{A}_{\mathcal{G}}$ is the adjacency matrix of $\mathcal{G}$. We use the notation $\phi_{\mathcal{G}}(t)$ for the characteristic polynomial of $\mathcal{G}$.
In [15] by Harary and in [23] by Sachs, the following theorem is proven, which calculates the coefficients of the characteristic polynomial $\mathcal{G}$ in terms of the counts of the elementary subgraphs of $\mathcal{G}$:
@@
Given a graph $\mathcal{G}$ of order $n$, then we have the following formula for the co-degree $d$-th coefficient of $\phi_{\mathcal{G}}(t)$:
@@
    [${\mathcal{G}}{X}$ ] is the number of isomorphic copies of $X$ in $\mathcal{G}$
@@
Next, we apply Theorem 2.8 and obtain equations relating number of walks and sums of heaps. Let $\mathcal{G}$ be a simple graph. For each cycle $\beta\in\widetilde{\mathcal{B}}(\mathcal{G})$, we let $w(\beta):=t^{-|E(\beta)|}$, where $|E(\beta)|$ is the length of $\beta$. Recall that the weight of a heap is defined as the product of the weights of its pieces.
@@
Elementary subgraphs of $\mathcal{G}$ are of the form $X=Y_{1}\sqcup\ldots\sqcup Y_{r}\sqcup Z_{1}\sqcup\ldots\sqcup Z_{s}\in\mathcal{E}(\mathcal{G})$, where $\{Y_{i}\}_{i=1}^{r}$ are edges and $\{Z_{j}\}_{j=1}^{s}$ are cycles of length $\geq 3$. By orienting the cycles in two directions, we can make $2^{s}$ many trivial heaps from such an elementary subgraph. Therefore, the following equation holds:
$\widetilde{\phi}_{\mathcal{G}}(t)=\sum_{T\in\mathbbm{t}(\mathcal{G})}(-1)^{c(T)}w(T)=\sum_{X\in\mathcal{E}(\mathcal{G})}t^{-|E(X)|}(-1)^{c(X)}2^{z(X)}$
@@
$-\log(\widetilde{\phi}_{\mathcal{G}}(t))$ $=-\log\left(\sum_{X\in\mathbbm{t}(\mathcal{B})}(-1)^{c(X)}w(X)\right)=\sum_{P\in\mathbbm{p}(\mathcal{G})}\frac{1}{|P|}w(P)=\sum_{P\in\mathbbm{p}(\mathcal{G})}\frac{1}{|P|}t^{-|E(P)|}$
@@
$-\log(\widetilde{\phi}_{\mathcal{G}}(t))=\sum_{d\geq 1}\dfrac{1}{d}w_{d}t^{-d}=\sum_{\mathbf{w}\in\mathcal{U}(\mathcal{G})}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}$
@@
$-\log(\widetilde{\phi}_{\mathcal{G}}(t))=\sum_{X\in\textup{Inf}^{1}(\mathcal{H})}t^{-|E(X)|}\dfrac{|\mathfrak{C}(X)|}{M_{X}}$
@@
$=\sum_{X\in\textup{Inf}^{1}(\mathcal{G})}t^{-|E(X)|}\dfrac{1}{|E(X)|}\dfrac{|\mathcal{W}(X)|}{M_{X}}$
@@
$=\sum_{d\geq 0}t^{-d}\dfrac{1}{d}\sum_{X\in\textup{Inf}^{1}_{d}(\mathcal{G})}\dfrac{|\mathcal{W}(X)|}{M_{X}}=\sum_{\mathbf{w}\in\mathcal{U}(\mathcal{G})}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}$
where the last equality follows from the bijective correspondence $\mathcal{U}(\mathcal{G})\leftrightarrow\bigsqcup_{X\in\textup{Inf}^{1}(\mathcal{G})}\mathcal{W}(X)/\mathord{\sim}_{\mathcal{T}(X)}$, shown in the proof of Corollary 2.7, where each equivalence class $[w]_{\mathord{\sim}_{\mathcal{T}(X)}}\in\bigsqcup_{\textup{Inf}^{1}(\mathcal{G})}\mathcal{W}(X)/\mathord{\sim}_{\mathcal{T}(X)}$ has size $M_{X}$. ∎
@@
$\frac{1}{\widetilde{\phi}_{\mathcal{G}}(t)}=\frac{1}{\sum_{T\in\mathbbm{t}(\mathcal{G})}(-1)^{c(T)}w(T)}=\sum_{H\in\mathbbm{h}(\mathcal{G})}w(H)=\sum_{H\in\mathbbm{h}(\mathcal{G})}t^{-|E(H)|}$
@@
A classic theorem in algebraic graph theory measures the number of closed walks of a simple graph $\mathcal{G}$ at a vertex $u$ via the characteristic polynomial. In particular, it is known that ([28, p. 36] [26, 3]) the Heaps of Pieces framework yields a proof of this theorem, stated below. (In the next section, we use the same approach with hypergraphs.)
@@
Let $\mathcal{G}$ be a simple graph of rank $k=2$. Let $u\in V(\mathcal{G})$ be a vertex. Then,
$\dfrac{\widetilde{\phi}_{\mathcal{G}-u}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}=\sum_{d\geq 0}w^{u}_{d}t^{-d}$
In other words, the number of closed walks of $\mathcal{G}$ at $u$ of length $d$ is equal to the coefficient of $t^{-d}$ in the fraction $\dfrac{\widetilde{\phi}_{\mathcal{G}-u}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}$.
@@
Let ~$\widetilde{\mathcal{B}}^{u}(\mathcal{G})$, or shortly ~$\widetilde{\mathcal{B}}^{u}$ be the union of cycles and edgegons of $\mathcal{G}$, containing $u$. By Theorem 2.8, we have:
$\dfrac{\widetilde{\phi}_{\mathcal{G}-u}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}=\dfrac{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}}\setminus\widetilde{\mathcal{B}}^{u})}(-1)^{c(T)}w(T)}{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}})}(-1)^{c(T)}w(T)}=\sum_{\begin{subarray}{c}H\in\mathbbm{h}(\widetilde{\mathcal{B}})\\ \mathcal{M}(H)\subseteq\widetilde{\mathcal{B}}^{u}\end{subarray}}w(H)=\sum_{\begin{subarray}{c}P\in\mathbbm{p}(\widetilde{\mathcal{B}})\\ \mathcal{M}(P)\subseteq\widetilde{\mathcal{B}}^{u}\end{subarray}}w(P)$
@@
$=\sum_{\mathbf{w}\in\mathcal{U}^{u}(\mathcal{G})}t^{-|\mathbf{w}|}=\sum_{d\geq 0}w^{u}_{d}t^{-d}$ by Corollary 2.7 Part 2
@@
Similarly, for a simple graph $\mathcal{G}$ and an edge $e\in E(\mathcal{G})$, the number of walks ending at $e$ can be measured via the characteristic polynomial. Given a simple graph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$, then a simple graph $\mathcal{G}$ is a spanning subgraph of $\mathcal{H}$ provided $V(\mathcal{G})=V(\mathcal{H})$ and $E(\mathcal{G})\subseteq E(\mathcal{H})$ ([4, p. 46]) In particular, we write $\mathcal{H}-e$ for the subgraph obtained deleting $e$, without removing any vertex.
@@
Let $\mathcal{G}$ be a simple graph of rank $k=2$. Let $e\in E(\mathcal{G})$ be an edge. Let $\mathcal{G}-e$ be the spanning subgraph of $\mathcal{G}$, obtained by removing $e$. Then,
$\dfrac{\widetilde{\phi}_{\mathcal{G}-e}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}=\sum_{d\geq 0}w^{e}_{d}t^{-d}$
In other words, the number of closed walks of $\mathcal{G}$ ending at $e$ of length $d$ is equal to the coefficient of $t^{-d}$ in the fraction $\dfrac{\widetilde{\phi}_{\mathcal{G}-e}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}$.

### v3

@@ Abstract
It is a classical result due to Jacobi in algebraic combinatorics that the generating function of closed walks at a vertex $u$ in a graph $G$ is determined by the rational function
$\frac{\phi_{G-u}(t)}{\phi_{G}(t)}$
where $\phi_{G}(t)$ is the characteristic polynomial of $G$. In this paper, we show that the corresponding rational function for a hypergraph is also a generating function for some combinatorial objects in the hypergraph.
We make use of the Heaps of Pieces framework, developed by Viennot, demonstrating its use on graphs, digraphs, and multigraphs before using it on hypergraphs. In the case of a graph $G$, the pieces are cycles and the concurrence relation is sharing a vertex. The pyramids with maximal piece containing a vertex $u\in V(G)$ are in one-to-one correspondence with closed walks at $u$. In the case of a hypergraph $\mathcal{H}$, connected “infragraphs” can be defined as the set of pieces, with the same concurrence relation: sharing a vertex. Our main results are established by analyzing multivariate resultants of polynomial systems associated to adjacency hypermatrices.
@@ 1 Introduction
Given a simple graph $G$, let $\phi_{G}(t)$ be the characteristic polynomial of its adjacency matrix $\mathbb{A}_{G}$. For a vertex $u\in V(G)$, a well-known classical identity ([12, Equation (1), p. 52]) is,
$\sum_{d\geq 0}w_{d}^{u}t^{d}=\frac{t^{-1}\phi_{G-u}(t^{-1})}{\phi_{G}(t^{-1})}$ (1.1)
where the $d$-th coefficient $w_{d}^{u}$ is the number of closed walks of length $d$ at $u$, and $G-u$ is the vertex-deleted subgraph of $G$, obtained by deleting the vertex $u$, together with all the edges incident with $u$. This result is attributed to Jacobi in 1833 ([16], see also [12, 11]). Note that the rational function on the right-hand side is equal to the generating function on the left-hand side, with non-negative integer coefficients. Thus, the function in question is an example of a broader class of functions of interest, the rational generating functions, investigated in [5].
@@
The theory of Heaps of Pieces is developed by Viennot (see [28, p. 36], [26, 3, 5, 18]). In the case of a finite simple graph $G$, one takes cycles of $G$ as pieces, and sharing a vertex as the concurrence relation. In Section 2.4, this approach is demonstrated and and laid out in detail, both for completeness and for reference. In particular, Corollary 2.7 establishes the bijection between closed walks of $G$ at a fixed vertex $u$, and the pyramids with maximal piece containing $u$.
@@
When two infragraphs are disconnected, we show that the weight of the disjoint union is the product of the weights of the components, so the multiplicativity condition is satisfied. This setting generalizes the Heaps of Pieces framework outlined above for graphs of rank $2$ in the following sense: The weight of a connected infragraph $X$ of a graph $G$ of rank $2$ is non-zero only if $X$ is indecomposable, i.e., it is a cycle or an edge (see [8]).
@@ 2.2 Operations on the Trails of an Eulerian Digraph
We denote by $E(\mathbf{w})=\{e_{1},\ldots,e_{d}\}$ the set of edges and by $V(\mathbf{w})=\{v_{0},\ldots,v_{d}\}$ the set of vertices of a walk $\mathbf{w}$. A walk $\mathbf{w}$ is called a trail, if its edges are distinct. A closed trail $\mathbf{w}$ of $X$ is an Eulerian trail when $E(\mathbf{w})=E(X)$. A path is a trail without repeated vertices. A closed trail without repeated internal vertices is said to be a simple closed trail. We denote by $\mathcal{T}(X)$ and $\mathcal{W}(X)$, the sets of closed trails and Eulerian (closed) trails of $X$, respectively. Given a vertex $u\in V(X)$, we write $\mathcal{T}^{u}(X)$ for the set of closed trails of $X$ at $u$ and $\mathcal{W}^{u}(X)$ for the set of Eulerian trails of $X$ at $u$. Given an edge $e\in E(X)$, we denote by $\mathcal{T}^{e}(X)$, the set of closed trails of $X$ ending at $e$ and by $\mathcal{W}^{e}(X)$, the set of Eulerian trails of $X$ ending at $e$.
@@ 2.4 Bijections between Pyramids and Trails/Walks
where, $v_{m}=v_{0}^{\prime}$ and $0\leq m,r$. Then, the concatenation of $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, denoted $\mathbf{w}_{1}\mathbf{w}_{2}$, is the trail:
@@
Trails of a multi-graph do not visit the same edge more than once. Given a simple graph $G$, we have a bijection (cf. Corollary 2.7 below.) featuring walks in which the same edge may be visited more than once and pyramids made of not necessarily distinct pieces. Given a sequence of vertices of a simple graph $G$, there is at most one walk visiting the sequence of vertices in the given order. This means that a walk is uniquely determined by its sequence of vertices.
Let $G$ be a connected simple graph of rank $2$. As defined in Section 2.1, the set of cycles of $G$ is denoted $\mathcal{B}(G)$. Since the graph $G$ is simple, it can not have a digon (a cycle of length $2$). However, walks of a graph of length $2$ up to cyclic permutation of edges, called edgegons, are included in the set of pieces of $G$: Let $\mathcal{Q}(G):=\{[(v_{0},e,v_{1},e,v_{0})]_{\rho}:e=\{v_{0},v_{1}\}\in E(G)\}$. Write $\widetilde{B}(G):=\mathcal{B}(G)\sqcup\mathcal{Q}(G)$ for the union of the set of cycles (of length $\geq 3$) and the set of edgegons of $G$. (We have $[(v_{0},e,v_{1},e,v_{0})]_{\rho}=[(v_{1},e,v_{0},e,v_{1})]_{\rho}$, for each edge $e=\{v_{0},v_{1}\}\in E(G)$ and necessarily, a graph $G$ has $|E(G)|$ many edgegons.)
@@
    Consider the simple graph $G$, obtained by appending an edge to a triangle. We list the pieces of $G$ below.
@@
By Veblen’s Theorem ([25, p. 87]), a connected simple graph $G$ has an Eulerian trail, i.e., $\mathcal{W}(G)\neq\emptyset$ if and only if it is obtained by an edge-disjoint union of cycles of length $\geq 3$.
Next, we derive a bijection involving walks of a simple graph $G$ and pyramids made of cycles and edgegons.
@@
Let $G$ be a simple connected graph, $u\in V(G)$ and $e\in E(G)$. Define the mapping
$q:\mathcal{U}(G)\rightarrow\mathbbm{p}(G)$
@@
    $q\big{|}_{\mathcal{U}^{e}(G)}:\mathcal{U}^{e}(G)\rightarrow\mathbbm{p}^{e}(\widetilde{\mathcal{B}}(G))$ is a bijection between closed walks of $G$ ending at $e$ and pyramids of $G$, with maximal piece containing $e$.
@@
    $q\big{|}_{\mathcal{U}^{u}(G)}:\mathcal{U}^{u}(G)\rightarrow\mathbbm{p}^{u}(\widetilde{\mathcal{B}}(G))$ is a bijection between closed walks of $G$ at $u$ and pyramids of $G$, with maximal piece containing $u$.
@@ 2.5 The Characteristic Polynomial of a Graph as a sum over Trivial Heaps
We intend to apply Theorem 2.8 to the characteristic polynomial of a simple graph $G$ of order $n$ and rank $2$, which is defined as the determinant $\text{det}(t\cdot I_{n}-\mathbb{A}_{G})$, where $I_{n}$ is the identity matrix of dimension $n$ and $\mathbb{A}_{G}$ is the adjacency matrix of $G$. We use the notation $\phi_{G}(t)$ for the characteristic polynomial of $G$.
In [15] by Harary and in [23] by Sachs, the following theorem is proven, which calculates the coefficients of the characteristic polynomial $G$ in terms of the counts of the elementary subgraphs of $G$:
@@
Given a graph $G$ of order $n$, then we have the following formula for the co-degree $d$-th coefficient of $\phi_{G}(t)$:
@@
    [${G}{X}$ ] is the number of isomorphic copies of $X$ in $G$
@@
Next, we apply Theorem 2.8 and obtain equations relating number of walks and sums of heaps. Let $G$ be a simple graph. For each cycle $\beta\in\widetilde{\mathcal{B}}(G)$, we let $w(\beta):=t^{-|E(\beta)|}$, where $|E(\beta)|$ is the length of $\beta$. Recall that the weight of a heap is defined as the product of the weights of its pieces.
@@
Elementary subgraphs of $G$ are of the form $X=Y_{1}\sqcup\ldots\sqcup Y_{r}\sqcup Z_{1}\sqcup\ldots\sqcup Z_{s}\in\mathcal{E}(G)$, where $\{Y_{i}\}_{i=1}^{r}$ are edges and $\{Z_{j}\}_{j=1}^{s}$ are cycles of length $\geq 3$. By orienting the cycles in two directions, we can make $2^{s}$ many trivial heaps from such an elementary subgraph. Therefore, the following equation holds:
$\widetilde{\phi}_{G}(t)=\sum_{T\in\mathbbm{t}(G)}(-1)^{c(T)}w(T)=\sum_{X\in\mathcal{E}(G)}t^{-|E(X)|}(-1)^{c(X)}2^{z(X)}$
@@
$-\log(\widetilde{\phi}_{G}(t))$ $=-\log\left(\sum_{X\in\mathbbm{t}(\mathcal{B})}(-1)^{c(X)}w(X)\right)=\sum_{P\in\mathbbm{p}(G)}\frac{1}{|P|}w(P)=\sum_{P\in\mathbbm{p}(G)}\frac{1}{|P|}t^{-|E(P)|}$
@@
$-\log(\widetilde{\phi}_{G}(t))=\sum_{d\geq 1}\dfrac{1}{d}w_{d}t^{-d}=\sum_{\mathbf{w}\in\mathcal{U}(G)}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}$
@@
$-\log(\widetilde{\phi}_{G}(t))=\sum_{X\in\textup{Inf}^{1}(\mathcal{H})}t^{-|E(X)|}\dfrac{|\mathfrak{C}(X)|}{M_{X}}$
@@
$=\sum_{X\in\textup{Inf}^{1}(G)}t^{-|E(X)|}\dfrac{1}{|E(X)|}\dfrac{|\mathcal{W}(X)|}{M_{X}}$
@@
$=\sum_{d\geq 0}t^{-d}\dfrac{1}{d}\sum_{X\in\textup{Inf}^{1}_{d}(G)}\dfrac{|\mathcal{W}(X)|}{M_{X}}=\sum_{\mathbf{w}\in\mathcal{U}(G)}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}$
where the last equality follows from the bijective correspondence $\mathcal{U}(G)\leftrightarrow\bigsqcup_{X\in\textup{Inf}^{1}(G)}\mathcal{W}(X)/\mathord{\sim}_{\mathcal{T}(X)}$, shown in the proof of Corollary 2.7, where each equivalence class $[w]_{\mathord{\sim}_{\mathcal{T}(X)}}\in\bigsqcup_{\textup{Inf}^{1}(G)}\mathcal{W}(X)/\mathord{\sim}_{\mathcal{T}(X)}$ has size $M_{X}$. ∎
@@
$\frac{1}{\widetilde{\phi}_{G}(t)}=\frac{1}{\sum_{T\in\mathbbm{t}(G)}(-1)^{c(T)}w(T)}=\sum_{H\in\mathbbm{h}(G)}w(H)=\sum_{H\in\mathbbm{h}(G)}t^{-|E(H)|}$
@@
A classic theorem in algebraic graph theory measures the number of closed walks of a simple graph $G$ at a vertex $u$ via the characteristic polynomial. In particular, it is known that ([28, p. 36] [26, 3]) the Heaps of Pieces framework yields a proof of this theorem, stated below. (In the next section, we use the same approach with hypergraphs.)
@@
Let $G$ be a simple graph of rank $k=2$. Let $u\in V(G)$ be a vertex. Then,
$\dfrac{\widetilde{\phi}_{G-u}(t)}{\widetilde{\phi}_{G}(t)}=\sum_{d\geq 0}w^{u}_{d}t^{-d}$
In other words, the number of closed walks of $G$ at $u$ of length $d$ is equal to the coefficient of $t^{-d}$ in the fraction $\dfrac{\widetilde{\phi}_{G-u}(t)}{\widetilde{\phi}_{G}(t)}$.
@@
Let ~$\widetilde{\mathcal{B}}^{u}(G)$, or shortly ~$\widetilde{\mathcal{B}}^{u}$ be the union of cycles and edgegons of $G$, containing $u$. By Theorem 2.8, we have:
$\dfrac{\widetilde{\phi}_{G-u}(t)}{\widetilde{\phi}_{G}(t)}=\dfrac{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}}\setminus\widetilde{\mathcal{B}}^{u})}(-1)^{c(T)}w(T)}{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}})}(-1)^{c(T)}w(T)}=\sum_{\begin{subarray}{c}H\in\mathbbm{h}(\widetilde{\mathcal{B}})\\ \mathcal{M}(H)\subseteq\widetilde{\mathcal{B}}^{u}\end{subarray}}w(H)=\sum_{\begin{subarray}{c}P\in\mathbbm{p}(\widetilde{\mathcal{B}})\\ \mathcal{M}(P)\subseteq\widetilde{\mathcal{B}}^{u}\end{subarray}}w(P)$
@@
$=\sum_{\mathbf{w}\in\mathcal{U}^{u}(G)}t^{-|\mathbf{w}|}=\sum_{d\geq 0}w^{u}_{d}t^{-d}$ by Corollary 2.7 Part 2
@@
Similarly, for a simple graph $G$ and an edge $e\in E(G)$, the number of walks ending at $e$ can be measured via the characteristic polynomial. Given a simple graph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$, then a simple graph $G$ is a spanning subgraph of $\mathcal{H}$ provided $V(G)=V(\mathcal{H})$ and $E(G)\subseteq E(\mathcal{H})$ ([4, p. 46]) In particular, we write $\mathcal{H}-e$ for the subgraph obtained deleting $e$, without removing any vertex.
@@
Let $G$ be a simple graph of rank $k=2$. Let $e\in E(G)$ be an edge. Let $G-e$ be the spanning subgraph of $G$, obtained by removing $e$. Then,
$\dfrac{\widetilde{\phi}_{G-e}(t)}{\widetilde{\phi}_{G}(t)}=\sum_{d\geq 0}w^{e}_{d}t^{-d}$
In other words, the number of closed walks of $G$ ending at $e$ of length $d$ is equal to the coefficient of $t^{-d}$ in the fraction $\dfrac{\widetilde{\phi}_{G-e}(t)}{\widetilde{\phi}_{G}(t)}$.